OpenStudy (anonymous):

what is 2(x + 6) - (x + 1) = 16

OpenStudy (nottim):

isolate for x, really.

OpenStudy (anonymous):

2(5+6)-(5+1)=16 x=5

OpenStudy (nottim):

u want me to be more specific, or give you da answeR? like wat anon might do?

OpenStudy (anonymous):

ive gotten to 2x-(x+1)=4

OpenStudy (nottim):

ok, how does the negative sign outside a bracket, affect the values inside da bracket?

OpenStudy (anonymous):

im not for sure

OpenStudy (nottim):

it makes everything inside the opposite sign (positive to negative, negative to positive).

OpenStudy (anonymous):

what is -1/2x+2=-x+7

OpenStudy (nottim):

ERR.ISOLATE FOR X.

OpenStudy (nottim):

u good